What is the probability of rolling something other than double 4s using two six-sided dice? Round your answer to three decimal places.
Answer #1

Number of outcomes on roll of a die = 6

Number of outcomes when 2 six-sided dice are rolled = 6 x 6

Probability of rolling a 4 = 1/6

P(rolling double 4s) = 1/6 x 1/6 = 1/36

P(rolling something other than double 4s) = 1 - P(rolling double 4s)

= 1 - 1/36

= 35/36

= 0.972